Monthly Archives: July 2014

140723 LML

Kona arrived at 11 am.  The doors were open, no sample, the system had crashed and the red wire to the small fan was broken.  The strand frames were taken to the sink and washed with 409 and DI water, the bottle was rinsed with DI water, and the rest of the apparatus was washed in place with 409 and DI water.  During cleaning the system failed again, it was reset once more and activated.

140723 Pepperwood Observations

Celeste Dodge activated the collector today, hoping for a foggy night. The sample jar was still in place after the 7/22 cleaning and was left as found on 7/23. RH was 53 and Temp 29 at 3:20pm.

Also collected 53mL from Alicia Torregrosa’s passive juvik or ‘harp’ style collector, which was installed on 7/20 for comparison to the Shaumenner and ‘German’ style passives collectors installed at Pepperwood.

140722 Pepperwood Observations

On 7/22/14 Celeste Dodge collected the full fog sample and cleaned the collector, but opted to shutdown due to forecasts of high clouds. At 8:40am the RH was 94% and Temp was 15 deg C. It looks like the met station recorded over 100% RH at that time and the Temp was around 15.

The collector has been turning on very early (before it becomes foggy) and staying on until way after the fog lifts. We shut down as the fog lifted in the a.m. of 7/22 because we wanted to avoid an opening on the night of 7/22 when fog was unlikely.

140722 Fritzche Field

No sample collected, although doors were open.  Collector was rinsed and closed up.  Spider trap placed beneath old wind profiler setup.  Small leak noted between teflon and plexiglass.  May not be an issue, but rinse water droplet were coming out of it.

140723 ISB(UCSC)

Robert arrived at ISB at 8:30. The doors were open but no sample had been collected. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.

140721 UCSC

Kona arrived at 8 am.  The collector doors were open, no sample collected and the system had crashed.  DI water wash with everything in place; the bottle was rinsed at the sink.   During the wash the system failed again; Kona unplugged the main power, made sure all other connections were secure, then reset and activated the system.  Ready for collection.

140720 UCSC

No sample taken.  system crashed overnight due to communications error. RH sensor has checked out fine.  Trying to remotely log into system.  Strands cleaned with DI water in place.  Jar rinsed and reused.

140719 UCSC

No sample taken.  system crashed overnight due to communications error. RH sensor has checked out fine.  Trying to remotely log into system.  Strands cleaned with DI water in place.  Jar rinsed and reused.

140720 LML Observations

Isaac arrived at the collector at 9:50 am. The collector doors had not opened and no sample had been collected. Upon arrival the small fan was running. The system was then tested to ensure the program was operational. The test showed that the system was working as intended. The collector was then disassembled and its components were rinsed with DI water. The collector was then rinsed with DI water and the collector was then reassembled and reset for the next fog event.